OpenStudy (anonymous):

Please help me solve 1- 4(c-3) + 7 = 59 2-5g+4(5+3g) = 2 - g 3- k- 3k = 6k + 5 - 8k 4- -6a + 3 = -3(2a-1) Thanks

Theoceangirl (the_ocean_girl):

you need somebody to walk you through all 4 of them?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

:)

Theoceangirl (the_ocean_girl):

okay i'll take you through all of them so you understand

OpenStudy (anonymous):

thank you

Theoceangirl (the_ocean_girl):

no problem. okay do you know the first step for #1?

OpenStudy (anonymous):

divide. Maybe the 4 from each side?

OpenStudy (anonymous):

distribute

Theoceangirl (the_ocean_girl):

distribute it to what's in the parentheses first

Theoceangirl (the_ocean_girl):

then we do the reverse order of operations

OpenStudy (anonymous):

okay 4(c-3) = 4c - 12 + 7 = 59

Theoceangirl (the_ocean_girl):

now combine like terms

OpenStudy (anonymous):

4c - 12 + 7

Theoceangirl (the_ocean_girl):

-12 and +7 are like terms combine them and then add that to 59

OpenStudy (anonymous):

c = 54

Theoceangirl (the_ocean_girl):

you'll be left with 4c divide each side by 4 and that's how you get what c equals

OpenStudy (anonymous):

c = 13.5

Theoceangirl (the_ocean_girl):

you should have gotten -5 when you combine like terms to get rid of the -5 you have to add 5

Theoceangirl (the_ocean_girl):

sorry i should have said that 1st

OpenStudy (anonymous):

okay thats okay

Theoceangirl (the_ocean_girl):

it'll be 59+5 and then divide by 4

OpenStudy (anonymous):

14.75?

Theoceangirl (the_ocean_girl):

59+5= 64 64/4=16

OpenStudy (anonymous):

yay thank you

Theoceangirl (the_ocean_girl):

no problem! let's move on to #2

OpenStudy (anonymous):

oky does #2 equal 37

Theoceangirl (the_ocean_girl):

how did you get to that answer

OpenStudy (anonymous):

5g+4(5 + 3g) = 2 - g 5g + 4(8g) = 2-g 5g+32g = 2 - g 37g = 1g 37 = g

Theoceangirl (the_ocean_girl):

you can't combine 5 and 3g because 5 doesn't have a g that threw off your math you were supposed to begin with the distributive property again

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

5g + 20 + 12g = 2 - g

OpenStudy (anonymous):

then do I combine the like terms?

OpenStudy (anonymous):

hello?

Theoceangirl (the_ocean_girl):

sorry i was taking notes for a class

OpenStudy (anonymous):

oh thats cool

Theoceangirl (the_ocean_girl):

yes you combine like terms now so 5g+3g

OpenStudy (anonymous):

kay

Theoceangirl (the_ocean_girl):

**5g+12g

OpenStudy (anonymous):

add?

Theoceangirl (the_ocean_girl):

yes add 5g and 12g

OpenStudy (anonymous):

17g

Theoceangirl (the_ocean_girl):

okay so now you have 20+17g= 2-g

OpenStudy (anonymous):

okay Do I combine 17g and g or just add regularly?

OpenStudy (anonymous):

thanks Im sorry but I have to go

Theoceangirl (the_ocean_girl):

okay